OpenStudy (anonymous):

f(x) = x – 5. Write the function rule for g(x), which is f(x) translated two units upward. Type your answer as a function rule for g(x), like this: g(x) = 14x - 12

OpenStudy (anonymous):

f(x) = x – 5 g(x) = f(x) + 2 g(x) = x – 5 + 2 g(x) = x – 3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

add 2

OpenStudy (misty1212):

\[x-5+2=x-3\]